# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v3

@@ IV.3 Validity
V Numerical results for Floquet engineering
@@
V.1 Floquet Weyl nodes and total 3D Hall conductivity
Figure 6: (a) Floquet Weyl-like nodes [Eq. (28)] and (b) Floquet Weyl nodes [Eq. (34)] as functions of light intensity $A_{0}$. The parameters are: $\varphi\!=\!0$ (linearly polarized laser), $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
Based on Eqs. (28) and (34), the positions of the Floquet Weyl-like nodes and Floquet Weyl nodes can be plotted, as shown in Fig. 6. Figure 6(a) and (b) illustrate that the locations of these nodes are tunable via the light intensity $A_{0}$. Notably, when $A_{0}$ reaches a critical value of approximately $1.034$ (nm)-1, the two Floquet Weyl-like nodes (or Floquet Weyl nodes) merge into a single point, indicating a topological phase transition.
@@ V.1 Floquet Weyl nodes and total 3D Hall conductivity
V.2 Floquet Chern number
Figure 8: Chern numbers for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) correspond to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.
We compute the Chern numbers for both the Floquet Weyl-like model and the Floquet Weyl semimetal model, as shown in Fig. 8. The results indicate that the phase transition points, represented by the Floquet Weyl nodes, shift with varying light intensities. Detailed derivations of the Chern numbers for both Floquet models are provided in Sections SVIII and SIX in the Supplemental Material Sup , respectively.
V.3 Floquet energy spectrum
Figure 9: Energy spectra for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal with $k_{x}=k_{y}=0$. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines mark the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.
The energy spectra for the Floquet Weyl-like model and the Floquet Weyl semimetal, calculated at $k_{x}\!=\!k_{y}\!=\!0$, are shown in Fig. 9. The two green vertical dashed lines are used to label the positions of the two Weyl-like (or Weyl) points. The energy gapless points, corresponding to the Weyl nodes, exhibit a dependence on the light intensity.
V.4 Floquet Fermi arc
Figure 10: Energy spectra (solid curves) and Fermi arcs (horizontal line at $E\!=\!0$) for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal under open boundary conditions (OBCs) along the $y$ direction, with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.
In Fig. 10, we show the energy spectra (solid curves) and Fermi arcs (depicted as horizontal straight lines at $E\!=\!0$) for both the Floquet Weyl-like model and the Floquet Weyl semimetal under OBCs along the $y$-direction, with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. We observe that the lengths of the Floquet Fermi arcs change with light intensity, reflecting variations in the Floquet Weyl nodes.
VI Experimental Realizations
@@
Unlike the 3D cubic lattices employed in cold-atom proposals Dubček et al. (2015); Roy et al. (2018) and experimental realizations Wang et al. (2021); Lu et al. (2020), we propose a design based on a stack of 2D square lattices, with each plaquette in every principal plane hosting a flux of $\pi$, as described in Ref. Ningyuan et al. (2015). This configuration corresponds to a system of spin-$1/2$ particles with engineered spin-dependent tunneling.
@@
The system thus takes the form of a multi-layer 2D time-reversal symmetric tight-binding model, with two sites per unit cell. This differs from the 3D time-reversal symmetric tight-binding model on a cubic lattice, which was originally proposed in Ref. Lu et al. (2019) to realize Weyl semimetals in 3D momentum space.
@@
To implement this multi-layer 2D square lattice in an electrical circuit, each lattice site is represented by an inductor, while tunnel couplings are modeled by pairs of capacitors. Zero-phase tunnel couplings connect the positive ends of adjacent inductors, while their negative ends are similarly linked. A $\pi$-phase tunnel coupling capacitively links the positive end of one inductor to the negative end of a neighboring inductor, and vice versa.
@@
The stacking of the 2D square lattice layers is coupled with a spatially periodic potential, which can be induced by an alternating current Qin et al. (2024) or via the charge-density-wave mechanism Zhang et al. (2023); Qin et al. (2020). This approach offers a promising platform for simulating the Weyl-like model, enabling experimental exploration of its novel topological properties.
VII Summary

### v4

@@ V.3 Validity
VI Numerical results for Floquet engineering
@@
VI.1 Floquet Weyl nodes and total 3D Hall conductivity
Figure 12: (a) Floquet Weyl-like nodes [Eq. (40)] and (b) Floquet Weyl nodes [Eq. (46)] as functions of light intensity $A_{0}$. The parameters are $\varphi\!=\!0$ (linearly polarized laser), $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
Based on Eqs. (40) and (46), the positions of the Floquet Weyl-like nodes and Floquet Weyl nodes can be plotted, as shown in Fig. 12. Figures 12(a) and 12(b) illustrate that the locations of these nodes are tunable via the light intensity $A_{0}$. Notably, when $A_{0}$ reaches a critical value of $\sim 1.034$ (nm)-1, the two Floquet Weyl-like nodes (or Floquet Weyl nodes) merge into a single point, indicating a topological phase transition.
@@ VI.1 Floquet Weyl nodes and total 3D Hall conductivity
VI.2 Floquet Chern number
Figure 14: Chern numbers for (a1) and (b1) the Floquet Weyl-like model and (a2) and (b2) the Floquet Weyl semimetal. Panels (a1) and (a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1) and (b2) correspond to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 12.
We compute the Chern numbers for both the Floquet Weyl-like model and the Floquet Weyl semimetal model, as shown in Fig. 14. The results indicate that the phase transition points, represented by the Floquet Weyl nodes, shift with varying light intensities. Detailed derivations of the Chern numbers for both Floquet models are provided in Sections SX and SXI in the Supplemental Material Sup , respectively.
VI.3 Floquet energy bands
Figure 15: Energy bands for (a1) and (b1) the Floquet Weyl-like model and (a2) and (b2) the Floquet Weyl semimetal with $k_{x}\!=\!k_{y}\!=\!0$. Panels (a1) and (a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1) and (b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines mark the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 12.
The energy bands for the Floquet Weyl-like model and the Floquet Weyl semimetal, calculated at $k_{x}\!=\!k_{y}\!=\!0$, are shown in Fig. 15. The two green vertical dashed lines are used to label the positions of the two Weyl-like (or Weyl) points. The energy gapless points, corresponding to the Weyl nodes, exhibit a dependence on the light intensity.
VI.4 Floquet Fermi arc
Figure 16: Energy bands (solid curves) and Fermi arcs (horizontal line at E=0E\!=\!0) for (a1) and (b1) the Floquet Weyl-like model and (a2) and (b2) the Floquet Weyl semimetal under open boundary conditions (OBCs) along the $y$ direction, with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. Panels (a1) and (a2) correspond to $A_{0}\!=\!0.1$ (nm)-1 and panels (b1) and (b2) to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 12.
In Fig. 16, we show the energy bands (solid curves) and Fermi arcs (depicted as horizontal straight lines at E=0E\!=\!0) for both the Floquet Weyl-like model and the Floquet Weyl semimetal under OBCs along the $y$ direction, with $k_{x}\!=\!0$ and $L_{y}\!=\!31a$. We observe that the lengths of the Floquet Fermi arcs change with light intensity, reflecting variations in the Floquet Weyl nodes.
VII Experimental Realizations
@@
Unlike the 3D cubic lattices employed in cold-atom proposals Dubček et al. (2015); Roy et al. (2018) and experimental realizations Wang et al. (2021); Lu et al. (2020), we propose a design based on a stack of 2D square lattices, with each plaquette in every principal plane hosting a flux of $\pi$, as described in Ref. Ningyuan et al. (2015). This configuration corresponds to a system of spin-$\frac{1}{2}$ particles with engineered spin-dependent tunneling.
@@
The system thus takes the form of a multilayer 2D time-reversal symmetric tight-binding model, with two sites per unit cell. This differs from the 3D time-reversal symmetric tight-binding model on a cubic lattice, which was originally proposed in Ref. Lu et al. (2019) to realize Weyl semimetals in 3D momentum space.
@@
To implement this multilayer 2D square lattice in an electrical circuit, each lattice site is represented by an inductor, while tunnel couplings are modeled by pairs of capacitors. Zero-phase tunnel couplings connect the positive ends of adjacent inductors, while their negative ends are similarly linked. A $\pi$-phase tunnel coupling capacitively links the positive end of one inductor to the negative end of a neighboring inductor and vice versa.
@@
The stacking of the 2D square lattice layers is coupled with a spatially periodic potential, which can be induced by an alternating current Qin et al. (2024a) or via the charge-density-wave mechanism Zhang et al. (2023); Qin et al. (2020). This approach offers a promising platform for simulating the Weyl-like model, enabling experimental exploration of its topological properties.
VIII Realization for both type-I and type-II Weyl-like model behaviors
VIII.1 Continuous Model
To investigate the emergence of both type-I and II Weyl-like behaviors, we consider a tilted continuous Weyl-like model Hamiltonian in the $(k_{x},k_{y},z)$ space, given by
${\cal H}({\bf k}_{||},z)$ $\!=\!$ $\left\{-M_{0}\!+\!C_{z}(Q_{z}z)^{2}\!+\!C_{||}[(k_{x}a)^{2}\!+\!(k_{y}a)^{2}]\right\}\!\sigma_{0}$ (48)
$\!+\left\{M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!-\!J_{||}[(k_{x}a)^{2}\!+\!(k_{y}a)^{2}]\right\}\!\sigma_{z}$
$\!+v_{0}k_{x}a\sigma_{x}\!+\!v_{0}k_{y}a\sigma_{y},$
where $Q_{z}$ is the wave vector of the spatially periodic potential along the $z$ direction; ${\bf k}_{||}\!=\!(k_{x},k_{y})$; $a$ is the lattice constant; $\sigma_{0}$ is a $2\times 2$ identity matrix; $M_{0}$, $C_{z}$, $C_{||}$, $M_{z}$, $J_{z}$, $J_{||}$, and $v_{0}$ are model parameters. The corresponding eigenenergies are
$E_{\pm}({\bf k}_{||},z)$
$\!=\!$ $-M_{0}\!+\!C_{z}(Q_{z}z)^{2}\!+\!C_{||}[(k_{x}a)^{2}\!+\!(k_{y}a)^{2}]$
$\pm\sqrt{(v_{0}a)^{2}(k_{x}^{2}\!+\!k_{y}^{2})\!+\!\left[M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!-\!J_{||}a^{2}(k_{x}^{2}\!+\!k_{y}^{2})\right]^{2}},$
which indicates two Weyl-like points at $z_{c\pm}\!=\!\pm(1/Q_{z})\sqrt{M_{z}/J_{z}}$ when $k_{x}\!=\!k_{y}\!=\!0$.
Expanding Eq. (VIII.1) around the Weyl-like points $(0,0,z_{c\pm})$ Chen et al. (2018c); Qin et al. (2020), we obtain the linearized energy bands as
$E_{\pm}(z)$ $\!\approx\!$ $-M_{0}\!+\!C_{z}(Q_{z}z_{c\pm})^{2}\!-\!2C_{z}Q_{z}^{2}z_{c\pm}z$ (50)
$\!\pm\Big|M_{z}\!-\!J_{z}(Q_{z}z_{c\pm})^{2}\!+\!2J_{z}Q_{z}^{2}z_{c\pm}z\Big|$
$\!=\!$ $-M_{0}\!+\!C_{z}\frac{M_{z}}{J_{z}}\!\mp\!2C_{z}\sqrt{\frac{M_{z}}{J_{z}}}Q_{z}z\pm\Big|2\sqrt{M_{z}J_{z}}Q_{z}z\Big|$
$\!=\!$ $\mp 2C_{z}\sqrt{\frac{M_{z}}{J_{z}}}Q_{z}z\pm\Big|2\sqrt{M_{z}J_{z}}Q_{z}z\Big|,$
where we used $(z\!-\!z_{c\pm})^{2}\!\approx\!z_{c\pm}^{2}\!-\!2z_{c\pm}z$ and assumed $M_{0}\!=\!C_{z}M_{z}/J_{z}$ for simplicity. The first term $\mp 2C_{z}\sqrt{M_{z}/J_{z}}Q_{z}z$ characterizes the tilt of the energy bands, while the second term $\pm 2\sqrt{M_{z}J_{z}}Q_{z}|z|$ describes the band splitting. When the tilt dominates the splitting, i.e., $\Big|2C_{z}\sqrt{M_{z}/J_{z}}Q_{z}z\Big|\!>\!\Big|2\sqrt{M_{z}J_{z}}Q_{z}z\Big|$, the Weyl-like point becomes type II. We define a tilt ratio as Chen et al. (2018c); Soluyanov et al. (2015); Li et al. (2017)
$F_{z}\!=\!\frac{\Big|2C_{z}\sqrt{M_{z}/J_{z}}Q_{z}z\Big|}{\Big|2\sqrt{M_{z}J_{z}}Q_{z}z\Big|}\!=\!\frac{C_{z}}{J_{z}}.$ (51)
Thus, for $F_{z}\!<\!1$ (i.e., $C_{z}\!<\!J_{z}$), the tilt is weak, resulting in a type-I Weyl-like point [Fig. 17(a)]. Conversely, for $F_{z}\!>\!1$ (i.e., $C_{z}\!>\!J_{z}$), the tilt is strong and both bands radiate in the same direction, indicating a type-II Weyl-like point [Fig. 17(b)]. Therefore, by tuning $C_{z}$ while keeping other parameters fixed, both type-I and II Weyl-like behaviors can be realized.
Figure 17: Energy bands of the tilted Weyl-like model with $k_{x}\!=\!k_{y}\!=\!0$ for (a) type I ($C_{z}\!=\!0.25$ eV <\!<\! $J_{z}\!=\!0.5$ eV) and (b) type II ($C_{z}\!=\!1$ eV >\!>\! $J_{z}\!=\!0.5$ eV). Vertical dashed green lines indicate the Weyl-like points. Other parameters: $M_{z}\!=\!0.5$ eV, $M_{0}\!=\!C_{z}M_{z}/J_{z}$, $v_{0}\!=\!1$ eV, and a=1a\!=\!1 nm.
As a dual picture, the tilted continuous Weyl semimetal model is given by
${\cal H}_{\rm W}({\bf k})$ $\!=\!$ $\left\{-M_{0}\!+\!C_{z}(k_{z}a)^{2}\!+\!C_{||}[(k_{x}a)^{2}\!+\!(k_{y}a)^{2}]\right\}\!\sigma_{0}$ (52)
$\!+\left\{M_{z}\!-\!J_{z}(k_{z}a)^{2}\!-\!J_{||}[(k_{x}a)^{2}\!+\!(k_{y}a)^{2}]\right\}\!\sigma_{z}$
$\!+v_{0}k_{x}a\sigma_{x}\!+\!v_{0}k_{y}a\sigma_{y},$
where ${\bf k}\!=\!(k_{x},k_{y},k_{z})$, and we have replaced $Q_{z}z$ by $k_{z}a$ in Eq. (48). Similar criteria apply here, allowing one to tune between type-I and II Weyl semimetal phases by adjusting $C_{z}$ and $J_{z}$.
VIII.2 Berry curvature and 2D Hall conductance
For the tilted continuous Weyl-like model in Eq. (48), the layer Berry curvature is given by Kubo (1957); Kubo et al. (1957); Thouless et al. (1982); Shen (2017); Qin et al. (2023a, 2024b, 2024a, 2025)
${\cal F}_{xy}^{(\pm)}({\bf k}_{||},z)$
$\!=\!$ $-2\varepsilon_{xyz}\!\frac{{\rm Im}\!\left\langle\!\psi_{\pm}\!\left|\!\frac{\partial{\cal H}}{\partial k_{x}}\right|\!\psi_{\mp}\!\right\rangle\!\!\left\langle\!\psi_{\mp}\!\left|\!\frac{\partial\!{\cal H}}{\partial k_{y}}\!\right|\!\psi_{\pm}\!\right\rangle}{\left(E_{\pm}-E_{\mp}\right)^{2}}$
$\!=\!$ $\mp\frac{v_{0}^{2}a^{2}\left[M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!+\!J_{||}a^{2}k_{||}^{2}\right]}{2\left\{v_{0}^{2}a^{2}k_{||}^{2}\!+\!\left[M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!-\!J_{||}a^{2}k_{||}^{2}\right]^{2}\right\}^{3/2}},$
[…]
To obtain a quantized 2D Hall conductance, we restrict attention to the type-I regime ($C_{z}\!<\!J_{z}$), where the Fermi level $E_{F}\!=\!0$ lies in the band gap, as shown in Fig. 17(a).
The $z$-dependent 2D Hall conductance for each layer is Qi et al. (2006); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a); Chen et al. (2025)
$\sigma_{H(\pm)}^{2D}(z)$ $\!=\!$ $\frac{e^{2}}{h}\frac{1}{2\pi}\int{\cal F}_{xy}^{(\pm)}({\bf k}_{||},z)d^{2}{\bf k}_{||}$
$\!=\!$ $\pm\frac{e^{2}}{2h}\frac{1}{2\pi}\int_{0}^{2\pi}d\phi\int_{0}^{\infty}dk_{||}^{2}\frac{\partial\cos\theta}{\partial k_{||}^{2}}$
$\!=\!$ $\mp\frac{e^{2}}{2h}\Big[\text{sgn}(J_{||})\!+\!\text{sgn}(M_{z}\!-\!J_{z}(Q_{z}z)^{2})\Big]\!,$
where $\text{sgn}(x)$ denotes the sign function, which indicates the sign of the variable $x$, and
$\cos\theta\!=\!\frac{M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!-\!J_{||}a^{2}k_{||}^{2}}{\left\{v_{0}^{2}a^{2}k_{||}^{2}\!+\!\left[M_{z}\!-\!J_{z}(Q_{z}z)^{2}\!-\!J_{||}a^{2}k_{||}^{2}\right]^{2}\right\}^{1/2}}.$
Therefore, the layer Chern number is given by
$C_{\pm}(z)\!=\!\mp\frac{1}{2}\Big[\text{sgn}(J_{||})\!+\!\text{sgn}(M_{z}\!-\!J_{z}(Q_{z}z)^{2})\Big]\!.$ (56)
Similarly, for the tilted continuous Weyl semimetal model in Eq. (52), the Berry curvature reads
${\cal F}_{xy}^{(\pm)}({\bf k})\!=\!\mp\frac{v_{0}^{2}a^{2}\left[M_{z}\!-\!J_{z}a^{2}k_{z}^{2}\!+\!J_{||}a^{2}k_{||}^{2}\right]}{2\left\{v_{0}^{2}a^{2}k_{||}^{2}\!+\!\left[M_{z}\!-\!J_{z}a^{2}k_{z}^{2}\!-\!J_{||}a^{2}k_{||}^{2}\right]^{2}\right\}^{3/2}}.$
Furthermore, the corresponding 2D Hall conductance is given by
$\sigma_{H(\pm)}^{2D}(k_{z})$ $\!=\!$ $\frac{e^{2}}{h}\frac{1}{2\pi}\int{\cal F}_{xy}^{(\pm)}({\bf k}_{||},k_{z})d^{2}{\bf k}_{||}$
$\!=\!$ $\mp\frac{e^{2}}{2h}\Big[\text{sgn}(J_{||})\!+\!\text{sgn}(M_{z}\!-\!J_{z}(k_{z}a)^{2})\Big]\!,$
and the corresponding Chern number becomes
$C_{\pm}(k_{z})\!=\!\mp\frac{1}{2}\Big[\text{sgn}(J_{||})\!+\!\text{sgn}(M_{z}\!-\!J_{z}(k_{z}a)^{2})\Big]\!.$ (59)
VIII.3 3D Hall conductivity
The 3D Hall conductivity for the tilted type-I Weyl-like model in Eq. (48) is computed as
$\sigma_{H(\pm)}^{3D}$ $\!=\!$ $\sum_{j_{z}=-L_{z}/(2a)}^{L_{z}/(2a)}\frac{\sigma_{H(\pm)}^{2D}(z)}{L_{z}}\!=\!\frac{e^{2}}{h}\sum_{j_{z}=z_{c-}/a}^{z_{c+}/a}\frac{C_{\pm}(z)}{L_{z}}$ (60)
$\!\approx\!$ $\mp\frac{2e^{2}}{hL_{z}aQ_{z}}\sqrt{\frac{M_{z}}{J_{z}}}\!,$
where $j_{z}\!=\!z/a$ is the layer index, $L_{z}$ is the system length along the $z$ direction, and we used $z_{c\pm}\!=\!\pm(1/Q_{z})\sqrt{M_{z}/J_{z}}$, and the layer Chern number $C_{\pm}(z)$ [Eq. (56)] satisfies
$C_{\pm}(z)$ $\!=\!$ $\!\left\{\begin{array}[]{l}\mp 1,~~z_{c-}\!<\!z\!<\!z_{c+}~{\rm and}~J_{||}\!>\!0;\\ 0,~~z\!<\!z_{c-}~{\rm or}~z\!>\!z_{c+}~{\rm and}~J_{||}\!>\!0.\end{array}\right$. (63)
Similarly, the 3D Hall conductivity of the type-II Weyl-like model also depends on $L_{z}$, owing to the summation over the layer index $j_{z}\!=\!z/a$ in its definition for the Weyl-like model.
For the tilted continuous type-I Weyl semimetal model in Eq. (52), the 3D Hall conductivity is given by
$\sigma_{H(\pm)}^{3D}$ $\!=\!$ $\int_{-\pi/a}^{\pi/a}\frac{dk_{z}}{2\pi}\sigma_{H(\pm)}^{2D}(k_{z})\!=\!\frac{e^{2}}{h}\int_{k_{c-}}^{k_{c+}}C_{\pm}(k_{z})\frac{dk_{z}}{2\pi}$ (64)
$\!=\!$ $\mp\frac{e^{2}}{ha\pi}\sqrt{\frac{M_{z}}{J_{z}}},$
where we used $k_{c\pm}\!=\!\pm(1/a)\sqrt{M_{z}/J_{z}}$, and the Chern number $C_{\pm}(k_{z})$ [Eq. (59)] satisfies
$C_{\pm}(k_{z})$ $\!=\!$ $\!\left\{\begin{array}[]{l}\mp 1,~~k_{c-}\!<\!k_{z}\!<\!k_{c+}~{\rm and}~J_{||}\!>\!0;\\ 0,~~k_{z}\!<\!k_{c-}~{\rm or}~k_{z}\!>\!k_{c+}~{\rm and}~J_{||}\!>\!0.\end{array}\right$. (67)
By comparing Eqs. (60) and (64), we find that the 3D Hall conductivity of the Weyl-like model explicitly depends on the system size $L_{z}$, while that of the Weyl semimetal is independent of system size. This finite-size dependence provides a clear and unique signature of mixed-space topology in the Weyl-like model.
IX Summary